# Poincaré inequality and Quantitative De Giorgi method for hypoelliptic operators

### v2

@@ 4. Proof of the Harnack inequality
Let the operators ${\mathcal{A}},{\mathcal{B}}$ be such that ${\mathcal{B}}$ is given by (H) and ${\mathcal{A}}$ is more specifically given by
@@
Let $f$ be a sub-solution to (1.1) on $Q_{1}$ with $0\leq f$ in $Q_{1}$ (local case) or on ${\mathbb{R}}^{d}\times B_{1}^{\otimes\kappa}\times[-1,0]$ with $0\leq f\leq 1$ (non-local case). Then there exist $\varepsilon>0$ and $r\in(0,1)$ so that $\int_{Q_{1}}f^{2}\leq\varepsilon$ implies $f\leq 1/2$ on $Q_{r}$.
Note that in the local case, a stronger statement holds where the $L^{\infty}$ norm of $f$ on $Q_{r}$ is controlled by a constant (depending on ${\mathcal{A}}$, ${\mathcal{B}}$ and $r$) times the $L^{2}$ norm of $f$ on $Q_{1}$. In the non-local case, variants of this stronger statement exist if $f$ is not only a sub-solution but also a super-solution; we refer to [12, 14]. The proof is based on an energy estimate and the gain of integrability of the fundamental solution of the Kolmogorov equation, local or non-local. This is classical and we refer to [8, 13, 2, 19, 10, 1]. The non-local case with more than one commutator is obtained by scaling arguments or by calculating the fundamental solutions in Fourier.
The second De Giorgi Lemma, also known as the intermediate value Theorem, writes
@@
Let the operators ${\mathcal{A}},{\mathcal{B}}$ be such that ${\mathcal{B}}$ is given by (H) and ${\mathcal{A}}$ is more specifically given by
@@
Let $f$ be a sub-solution to (1.1) with $f\leq 1$ on $Q_{1}$ (local case) or $0\leq f\leq 1$ on ${\mathbb{R}}^{d}\times B_{1}^{\otimes\kappa}\times[-3,0]$ (non-local case). Let $\delta_{1},\delta_{2}\in(0,1)$. Then there are constants $r_{0}>0$ and $\nu\in(0,1)$ and $\theta\in(0,1)$ such that whenever
@@
The proof that this lemma follows from our Poincaré inequality and the first lemma of De Giorgi by using the argument of [8, Subsection 3.2] in the local case (in fact the proof is simpler since we do not have the error terms in the Poincaré inequality that were present in [8]), and it follows from [13, Subsection 5.2] in the non-local case (note that we have removed the mollifications $F_{i}$ of the cutoff functions in the statement of [13] since they can be used in the proof but removed from the statement straightforwardly by making the border region small enough). These proofs are done in the case of one commutator but they extend straightforwardly to the case of several commutators because they do not depend on the transport part ${\mathcal{T}}=\partial_{t}+{\mathcal{B}}$ of the equation, but only on ${\mathcal{A}}$ and the Poincaré inequality and first lemma of De Giorgi.
With the first and second De Giorgi lemma at hand, one can deduce the “measure-to-pointwise estimate” and the weak Harnack inequality by following [8] in the local case, and [13] in the non-local case. These proofs are independent of the equation and only depend on the previously established functional inequalities. The only minor change is that the kinetic cylinders have a slightly more complicated scaling on the variables $(x^{(\kappa)},\dots,x^{(0)},t)$ but the technical changes needed are, although slightly tedious, clear.

### v3

@@ 4. Proof of the weak Harnack inequality
Let the operators $\mathcal{A}$$\mathcal{B}$} be such that $\mathcal{B}$} is given by (H) and $\mathcal{A}$} is more specifically given by
@@
Let $f$ be a sub-solution to (1.1) in $Q_{1}$, such that in the local case $0\leq f$ in $Q_{1}$, and such that in the non-local case, $0\leq f\leq 1$ in $B_{1}^{\otimes\kappa}\times{\mathbb{R}}^{d}\times[-1,0]$. Then there exist $\varepsilon>0$ and $r\in(0,1)$ so that $\int_{Q_{1}}f^{2}\leq\varepsilon$ implies $f\leq 1/2$ on $Q_{r}$.
In the local case, a stronger statement holds where the $L^{\infty}$ norm of $f$ on $Q_{r}$ is controlled by a constant (depending on $\mathcal{A}$}, $\mathcal{B}$} and $r$) times the $L^{2}$ norm of $f$ on $Q_{1}$. In the non-local case, variants of this stronger statement exist if $f$ is not only a sub-solution but also a super-solution; we refer to [13, 15]. The proof is based on an energy estimate and the gain of integrability of the fundamental solution of the Kolmogorov equation, local or non-local. This is classical and we refer to [8, 14, 2, 20, 11, 1]. The non-local case with $\kappa\geq 2$ is obtained by scaling arguments, or by calculating the fundamental solutions in Fourier.
As a next step, one needs to prove the second De Giorgi Lemma, also known as the intermediate value Theorem, which controls the gradient of a sub-solution by getting an explicit bound on the measures of the sets where the solution is in between two values. It reads as follows:
@@
Let the operators $\mathcal{A}$$\mathcal{B}$} be such that $\mathcal{B}$} is given by (H) and $\mathcal{A}$} is more specifically given by
@@
Let $0\leq f\leq 1$ be a sub-solution to (1.1) in $Q_{1}$. Let $\delta_{1},\delta_{2}\in(0,1)$. Then there are constants $r_{0}>0$ and $\nu\in(0,1)$ and $\theta\in(0,1)$ such that whenever
@@
The proof of this lemma follows from our Poincaré inequality and the first lemma of De Giorgi by using the argument of [8, Subsection 3.2] in the local case (in fact the proof is simpler since we do not have the error terms in the Poincaré inequality that were present in [8]), and it follows from [14, Subsection 5.2] in the non-local case (note that we have removed the mollifications $F_{i}$ of the cutoff functions in the statement of [14] since they can be used in the proof but removed from the statement by making the border region small enough). These proofs are done in the case of one commutator but they extend straightforwardly to the case of several commutators because they do not depend on the transport part ${\mathcal{T}}=\partial_{t}+{\mathcal{B}}$ of the equation, but only on $\mathcal{A}$} and the Poincaré inequality and first lemma of De Giorgi.
With the first and second De Giorgi lemma at hand, one can deduce the “measure-to-pointwise estimate” and the weak Harnack inequality by following [8] in the local case, and [14] in the non-local case, see [1] also. These proofs are independent of the equation and only depend on the previously established functional inequalities. The only minor change is that the kinetic cylinders have a slightly more complicated scaling on the variables $(x^{(\kappa)},\dots,x^{(0)},t)$ but the technical changes needed are, although slightly tedious, clear.